OpenStudy (anonymous):

What is the equation of a line parallel to the line y = 4x + 7 and passes through the point (3, 0)?

OpenStudy (anonymous):

y=4(x-3)

OpenStudy (123456789):

you know that the slope has to be the same, since the lines are parallel, so \[y= 4x + b\] you have the coordinates (3,0) so you can substitute them in the equation above to get \[ 0 = 4(3) +b\] so b = -12 substitute into the original equation and you get\[y= 4x -12\]